# Evaluating Sound Similarity Metrics for Differentiable, Iterative Sound-Matching

### v1

@@ IV-C The Synthesizers
Once the programs are transpiled to differentiable synthesizers, the iterative sound-matching procedure is:
  1. 1.
  2. 2.
  3. 3.
    Calculating the loss between the target and output
  4. 4.
    Applying gradient updates to the synthesizer parameters
  5. 5.
In the following subsections, we describe the programs and their parameters in more detail, followed by the methodology in Section IV-A and results in Section IV-D.
@@ IV-C4 SineSaw-AM
IV-D Post hoc Analysis of Loss Performance
As summarized in Figure 2, for each program and evaluation (or scoring) method, there are four sets 1000 bootstrapped values, corresponding to the four loss functions. A post-hoc test is conducted in two stages in order to determine whether there is a difference in performance between loss functions for each program, and if so, which are the best performers. The first stage determines whether there is a difference in group means, with the null hypothesis being that all groups have similar mean ranks. The second stage ranks the loss functions from best to worst using the non-parametric Scott-Knott test (NPSK) [58, 59]. The Kruskal-Wallis test is used for the first stage [70]; this test pools and ranks all evaluation measures for a program, then tests whether the differences in mean rank for all loss function groups is zero. The Kruskal-Wallis test calculates an H statistic that is compared to a chi-square distribution with k-1 degrees of freedom (k is the number of groups, or 4, and degrees of freedom is 3). Using the significance level of 0.05, if the H statistic is greater than the critical value of the chi-square distribution then the null hypothesis is rejected, meaning that at least one group has a mean rank significantly above others [70]. Table II shows the Kruskal-Wallis results for every program and evaluation method, and shows that significant differences between loss function performance measures (i.e., the bootstrapped distributions of the evaluation for each loss function) exist in all cases except MSS for the Noise-AM program.
@@ IV-D Post hoc Analysis of Loss Performance
IV-D1 Best Performers For Each Program
Based on the bootstrapped distribution of scores, the NPSK algorithm ranks the loss functions from 1 (best) to a maximum of 4 (worst). In cases where the distributions are similar, multiple loss functions can be clustered into the same rank. In Figure 3, we use color-coded violin plots to visualize the best performers for each program. The corresponding colors for each rank are 1 2 3 4
@@ IV-D1 Best Performers For Each Program
Figure 3: Distributions and ranks of the loss functions based on three different performance measures. From left to right, the performance measures are: Likert, P-Loss, and MSS. Higher values indicate better performance. Rank colors are 1 2 3 4
IV-D2 BP-Noise
@@ IV-D3 Add-SineSaw
IV-D4 Noise-AM
IP-Loss and manual hearing tests yield identical rankings, and MSS results vary for for ranks 2 to 4. It is not surprising that MSS results differ from the hearing tests, as we saw in Table II, MSS showed no significant differences between groups. As shown in Figure 3c, all models selected DTW_Envelope as the best performer. This may be due to DTW_Envelope’s focus on periodic changes in loudness.
IV-D5 SineSaw-AM
@@
IV-E Consistency In Rankings
IV-E1 Manual Ranks
@@
IV-F Takeaways
@@
An important takeaway is that the loss function that yields the best sound-matching outcomes varies depending on the synthesizer program. This is perhaps due to the interaction between how the parameters of the synthesizer influence the sound, and the core sonic features of the loss function.
We also see that the DTW_Envelope and SIMSE_Spec (to our knowledge, not previously used in iterative sound-matching), lead to the best results in three of the four programs. This shows room for further creative approaches to differentiable sound-similarity measures, particularly since we observed that SOTA loss functions can be ineffective depending on the design of a synthesizer.
@@
Finally, we surprisingly see somewhat consistent results between the rankings assigned by manual hearing tests (which we take as the ground truth), and automatic measures of P-Loss and MSS. Due to the simplicity of our programs and the occasional deviations, we cannot conclude that these automatic measures are substitutes for human hearing tests.
@@ V Summary, Weaknesses, and Conclusion
The main hypothesis tested here is whether the performance of differentiable loss functions is influenced by the synthesis techniques used for sound-matching. We created iterative sound-matching experiments by combining four different loss functions with four different sound synthesis programs. We ranked the performance of the iterative sound-matching pipelines for every loss function and program, and observed that the success of the pipeline (that is, how closely the output sound matches the target sound) is program dependent. In other words, different synthesizer programs work best with different loss functions. Notably, we see that our novel use of DTW and SIMSE based differentiable loss functions (DTW_Envelope and SIMSE_Spec) can outperform what are regarded as the SOTA loss functions in 3 of 4 cases, although this is highly synthesizer dependent.
P-Loss and MSS have frequently been used as automatic performance measures, yet their “preference” has rarely been compared to human rankings. We observed that automatic performance measures and manual listening tests were often in agreement, despite this, manual verification of sound-matching results remains a necessity for future works.
Weaknesses: While we cannot prove that a universally best similarity measure does not exist, we can advocate for more creativity in the field based on our findings. Compared to previous work, we presented a more cohesive approach to iterative sound-matching which utilizes a variety of loss functions and synthesis methods. However, there are many other methods of synthesis and sound-similarity that can be combined in practically infinite ways. Due to this large search space, we set arbitrary parameters for the various signal processing functions. We used bare-bones versions of STFT and JTFS with fixed parameters. We did not test complex synthesizers using parallel and sequential DSP functions. Arbitrary hyperparameters such as learning rate and max number of iterations were selected for the DL pipeline, and only the RMSProp optimizer was tested.
@@
Like nearly all previous work in sound-matching, the main measure of success here was the accurate replication of sounds (or synth parameters) rather than imitation of sounds outside of the synthesizer’s domain. Replication of arbitrary features of sounds is an important component of sound-design, though it is much harder to define or measure. Future works could explore loss functions which only measure certain characteristics of sound (such as DTW_Envelope), and whether they can pave the way for better imitation sound-matching.

### v2

@@ III-B4 SineSaw-AM
III-C Evaluation Methods
The two automatic evaluation methods used here are P-Loss and MSS, as described in Section II-C. For P-Loss, the parameters are normalized between 0-1 based on the valid ranges defined in the Faust program, and the L1 distance between the normalized parameters is calculated. MSS is computed using a hop size of $100$ samples, and FFT window lengths of $(512,1024,2048,4096)$.
Automatic sound evaluation methods are often unreliable, as their alignment with the subjective judgments of human listeners is uncertain. Therefore, listening tests are conducted by randomly sampling 40 experiment results for each program and loss function pair. Two of the authors then assign blinded similarity scores to the outputs and targets using a 5-point Likert scale (1 = no similarity, 5 = near identical) [60]. Using Spearman’s rank correlation [61, 62], we found a very strong rate of agreement between the human listeners (see Section III-G), a good indicator that the scores assigned by manual listeners can be treated as the ground truth. We then combine the ranks assigned by both authors, giving us 80 ranks per program and loss function pair.
III-D Training Loop and Gradient Calculations
Given a differentiable loss and synthesizer, the iterative sound-matching procedure is as follows:
  1. 1.
  2. 2.
  3. 3.
    Calculating the loss between the target and output
  4. 4.
    Applying gradient updates to the synthesizer parameters
  5. 5.
III-E Ranking Loss Functions
Figure 2 is a visualized summary of how the loss functions are ranked. We have four loss functions and four synthesizer programs. The maximum number of iterations set to 200, and 300 experiments for each loss/synthesizer combination. Each synthesizer program is paired with four loss functions and 300 experiments are conducted and evaluated automatically with P-Loss and MSS, and manually with Likert scores. With this approach, for each program and loss function pair, there are three distributions that can be used to rank the loss functions. Two distributions with 300 automatically assigned similarities (P-Loss and MSS) and one distribution with 80 Likert scores (combining the 40 ranks assigned by each author).
For consistency and statistical robustness, the distributions are upsampled to 1000 values using bootstrapping [64]. Bootstrapping gives an estimation of the distribution for the mean performance of each experiment. $k$ (set to 1000) samples of $n$ values (set to 100% of the values) are taken with replacement from the empirical distribution (list of 300 performance values for MSS and P-Loss or 80 for manual rankings). The mean performance is calculated for each of the $k$ samples. The $k$ estimates of the mean performance give us a bootstrapped distribution for each loss function.
[…]
A post-hoc test is conducted in two stages in order to determine whether there is a difference in performance between loss functions for each program, and if so, which are the best performers. The first stage determines whether there is a difference in group means, with the null hypothesis being that all groups have similar mean ranks. The second stage ranks the loss functions from best to worst using the non-parametric Scott-Knott test (NPSK) [65, 66]. The Kruskal-Wallis test is used for the first stage [67]; this test pools and ranks all evaluation measures for a program, then tests whether the differences in mean rank for all loss function groups is zero. The Kruskal-Wallis test calculates an H statistic that is compared to a chi-square distribution with k-1 degrees of freedom (k is the number of groups, or 4, and degrees of freedom is 3). Using the significance level of 0.05, if the H statistic is greater than the critical value of the chi-square distribution then the null hypothesis is rejected, meaning that at least one group has a mean rank significantly above others [67].
@@ III-E Ranking Loss Functions
III-F Best Performers For Each Program
Table II shows the Kruskal-Wallis results for every program and evaluation method, we see significant differences between loss function performance measures (i.e., the bootstrapped distributions of the evaluation for each loss function) in 11 of the 12 cases, with the exception of MSS for the Noise-AM program.
Based on the bootstrapped distribution of the scores, the NPSK algorithm ranks the loss functions from 1 (best) to a maximum of 4 (worst). In cases where the distributions are similar, multiple loss functions can be clustered into the same rank.
In Figure 3, we use color-coded violin plots to visualize the best performers for each program. The corresponding colors for each rank are 1 2 3 4.
@@ III-F Best Performers For Each Program
Figure 3: Distributions and ranks of the loss functions based on three different performance measures. From left to right, the performance measures are: MSS, P-Loss, and Likert. Higher values indicate better performance. Rank colors are 1 2 3 4.
III-F1 BP-Noise
@@ III-F2 Add-SineSaw
III-F3 Noise-AM
P-Loss and manual hearing tests yield identical rankings, and MSS results vary for ranks 2 to 4. It is not surprising that MSS results differ from the hearing tests, as we saw in Table II, MSS showed no significant differences between groups. As shown in Figure 3c, all models selected DTW_Envelope as the best performer. This may be due to DTW_Envelope’s focus on periodic changes in loudness.
III-F4 SineSaw-AM
@@
III-G Consistency In Rankings
III-G1 Manual Ranks
@@
The experiment results provide answers regarding the main hypothesis of this work as well as the secondary questions. In the following sections, we will discuss our key findings, caveats, and make recommendations for future research.
IV Discussion
IV-A Key Findings
@@
The most important takeaway regarding our main hypothesis is that the loss function that yields the best sound-matching outcomes varies depending on the synthesizer program. This is likely due to the interaction between how the parameters of the synthesizer influence the sound, and the core sonic features used by the loss function.
Q1: To what extent do automatic evaluation metrics agree with manual listening tests?
@@
We see somewhat consistent results between the rankings assigned by manual hearing tests (which we take as the ground truth), and automatic measures of P-Loss and MSS. With the exception of P-Loss in BP-Noise narrowly selecting a different spectrogram based loss, all measures of performance selected the same top performer. However, given the simplicity of our programs and the occasional deviations, we cannot conclude that these automatic measures are substitutes for human hearing tests. As many previous works have done (see Table I), the use of MSS and/or P-Loss as general loss functions does seem appropriate in general cases, however, the use of listening tests or custom made loss functions in specific contexts remains necessary.
Q2: Are DTW and SIMSE effective measures of loss? If so, in what contexts?
DTW_Envelope consistently outperformed other measures in amplitude-modulated synthesis, where envelope alignment is critical. SIMSE_Spec performed best in subtractive synthesis, where scale-invariance captures noise-filtering effects more robustly than L1-based measures. These results suggest that the advantages of loss functions are context-dependent, underscoring the need for further exploration of creative differentiable similarity measures.
Q3: Can iterative differentiable optimization be applied effectively with synthesizers using classical DSP functions?
The approach of designing DSP functions in Faust and transpiling to differentiable Faust code was convenient, yielded meaningful outputs, and enabled comparative evaluation of losses. We were able to run our 1200 experiments within 72 hours on a laptop without a GPU [^4] Lenovo ThinkPad T480, i5-8350U CPU at 1.70GHz, 32 GB RAM..
IV-B Practical Recommendations
Our findings not only provide answers to the research questions proposed in the introduction, but also serve as guidelines for practitioners in selecting appropriate similarity measures and synthesis methods for future experiments.
  • In lieu of listening tests, MSS or P-Loss are generally useful for large-scale benchmarking. However, we encourage confirmation of such results with manual listening, particularly if specificity is important.
  • For amplitude-modulated synthesis, DTW-based losses such as DTW_Envelope are recommended over spectrogram differences and JTFS.
  • For subtractive/noise-filtered synthesis, spectrogram based losses are most effective. Based on our results, SIMSE appears to be the better method of spectrogram comparisons relative to L1, but the amplitude agnosticism of SIMSE likely plays a role in its poor performance in additive synthesis.
  • Iterative differentiable optimization via the Faust-to-JAX pipeline is a viable strategy for defining various differentiable DSP functions, requiring only modest hardware resources and providing a more natural approach to synthesizer definition.
  • In our amplitude modulation experiments, JTFS did not outperform other measures, contradicting prior claims of its superiority for mesostructures [14]. Thus, our results call into question its effectiveness as a general mesostructure measure, and highlight a need for further research into its utility.
IV-C Illustrating Loss Landscapes
To complement our quantitative analysis, we provide examples of loss landscapes that illustrate why specific losses succeed or fail in certain synthesis contexts. These are illustrative aids rather than conclusive evidence, and highlight the importance of Loss Landscape Navigation discussed earlier in Section II-G.
We show the landscapes of the loss functions with simplified versions of BP-Noise (HP-Noise) and Noise-AM (S-Noise-AM), each reduced to a single parameter for one-dimensional visualization. HP-Noise uses only a high-pass filter with cut-off between 100–20,000 Hz, while S-Noise-AM varies only the modulation rate between 0.1–20 Hz. Loss values are normalized to 0–1 to allow comparison across functions.
Figure 5 shows the HP-Noise results. L1_Spec, SIMSE_Spec, and JTFS exhibit clear minima at the correct parameter (red dashed line), whereas DTW_Envelope remains relatively flat, making gradient-based optimization more difficult.
Figure 5 shows the S-Noise-AM results. Here, DTW_Envelope produces the smoothest and most informative landscape, while spectrogram losses are minimized near the target but lack consistent correlation with parameter distance. JTFS remains largely flat. This aligns with DTW_Envelope’s superior performance in amplitude-modulated synthesis observed in our experiments.
Figure 4: Loss landscapes for BP-Noise with only a high-pass filter parameter. L1_Spec, SIMSE_Spec, and JTFS show clear global minima near the correct parameter, while DTW_Envelope remains flat around the target.
Figure 5: Loss landscapes for a simplified Noise-AM synthesizer with only amplitude modulation parameter. DTW_Envelope exhibits the smoothest and most informative landscape, explaining its superior performance in amplitude-modulated synthesis.
IV-D Applicability to More Complex Synthesizers
The experiments in this study were conducted with differentiable synthesizers chosen to isolate fundamental synthesis principles, each with two parameters. While this level of simplicity facilitates controlled comparisons, it does not fully capture the richness of real-world synthesizers, which can have hundreds of parameters with various routes. As a result, the extent of the generalizability of our findings does require further research.
Nonetheless, we believe that the insights here can extend naturally to more complex domains. Depending on the feature of interest, the observed dependence of loss performance on synthesis method is likely to remain stable even with added complexities such as layers of modulation, filtering, or nonlinearity.
DTW_Envelope measures very specific features of audio, and its use in the recommended settings—where low frequency amplitude modulation matching is required—would likely yield the desired results regardless of synthesizer complexity. Likewise, the use of SIMSE_Spec or L1_Spec for matching filter cut-offs is likely to be successful regardless of the underlying sound.
@@ V Summary, Weaknesses, and Conclusion
The main hypothesis tested here is whether the performance of differentiable loss functions is influenced by the synthesis techniques used for sound-matching. We conducted systematic iterative sound-matching experiments by combining four different loss functions with four different sound synthesis programs. We ranked the performance of the iterative sound-matching pipelines for every loss function and program, and observed that the success of the pipeline (that is, how closely the output sound matches the target sound) is program dependent. In other words, different synthesizer programs work best with different loss functions. Notably, we see that our novel use of DTW and SIMSE based differentiable loss functions (DTW_Envelope and SIMSE_Spec) can outperform what are regarded as the SOTA loss functions in 3 of 4 cases, although their success is highly synthesizer dependent.
P-Loss and MSS have frequently been used as automatic performance measures, yet their “preference” has rarely been compared to human rankings. We observed that automatic performance measures and manual listening tests were generally in agreement, despite this, manual verification of sound-matching results remains a necessity in non-general experiments due to occasional divergences between automatic and manual tests.
Weaknesses: While we cannot prove that a universally best similarity measure does not exist, we can advocate for more creativity in the field. Compared to previous work, we presented a more cohesive approach to iterative sound-matching which utilizes a variety of loss functions and synthesis methods. However, there are many other methods of synthesis and sound-similarity that can be combined in practically infinite ways. Due to this large search space, we set arbitrary parameters for the various signal processing functions. We used bare-bones versions of STFT and JTFS with fixed parameters. We did not test complex synthesizers using parallel and sequential DSP functions. Arbitrary hyperparameters such as learning rate and max number of iterations were selected for the DL pipeline, and only the RMSProp optimizer was tested.
@@
Like nearly all previous work in sound-matching, the main measure of success here was the accurate replication of sounds (or synthesizer parameters) rather than imitation of sounds outside of the synthesizer’s domain. Replication of arbitrary features of sounds is an important component of sound-design, though it is much harder to define or measure. Future works could explore loss functions which only measure certain characteristics of sound (such as DTW_Envelope), and whether they can pave the way for better imitation in sound-matching.